Calculate the degree of dissociation of PCl5 if the vapour density at the given temperature is 70.

Grade:

1 Answers

Sunil Kumar FP
9 years ago

initial vapour densuty/vapour density at equilibrium=mole at equilibrium/mole initial

PCL5 -----PCL3 +CL2
let the initial concentration be C
concentration of PCL5=C-Cx,
x is the degree of dissociation
PCL3=Cx
CL2=Cx
total mole at equilibrium=C +Cx
initial mole=C therefore C +Cx/C =104.25/70
1+x=1.485
x=.485